### v1

# SOMONITOR: Explainable Marketing Data Processing and Analysis with Large Language Models

Abstract.
Online marketing faces formidable challenges in managing and interpreting immense volumes of data necessary for competitor analysis, content research, and strategic branding. It is impossible to review hundreds to thousands of transient online content items by hand, and partial analysis often leads to suboptimal outcomes and poorly performing campaigns. We introduce an explainable AI framework SOMONITOR that aims to synergize human intuition with AI-based efficiency, helping marketers across all stages of the marketing funnel, from strategic planning to content creation and campaign execution. SOMONITOR incorporates a CTR prediction and ranking model for advertising content and uses large language models (LLMs) to process high-performing competitor content, identifying core content pillars such as target audiences, customer needs, and product features. These pillars are then organized into broader categories, including communication themes and targeted customer personas. By integrating these insights with data from the brand’s own advertising campaigns, SOMONITOR constructs a narrative for addressing new customer personas and simultaneously generates detailed content briefs in the form of user stories that can be directly applied by marketing teams to streamline content production and campaign execution. The adoption of SOMONITOR in daily operations allows digital marketers to quickly parse through extensive datasets, offering actionable insights that significantly enhance campaign effectiveness and overall job satisfaction.
Digital Advertising, Ads Performance Prediction, Deep Learning, Large Language Model, Explainable AI
isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Multimedia and multimodal retrievalccs: Information systems Computational advertising
Figure 1. General pipeline of the SOMONITOR advertising analysis framework.
[…]
General pipeline of the SOMONITOR advertising analysis framework.
1. Introduction
The contemporary online marketing landscape is bursting with an overwhelming amount of content in the form of databases of advertisements, marketing campaigns of various competitors, properties of user audiences, and more. Marketers must not only sift through this information but also extract meaningful insights regarding audience segmentation, trending topics, and customer needs.
Recent advances in the field of large language models (LLMs) have brought them to the fore as the tools of choice in various domains (see Section 2). These models, trained on huge datasets, exhibit remarkable capabilities in both content analysis and content generation. In online marketing, modern LLMs are already able to offer a promising solution by processing and analyzing large datasets to identify key elements, facilitating a more targeted and effective marketing strategy. Moreover, they can help generate new advertisement content, both text-based and multimodal.
Despite all this progress, in this work we note that vanilla LLMs also have limitations in content scoring, which is a crucial aspect of effective marketing. We show that a system specifically tailored for CTR prediction based on advertising content performs better at content scoring than standard LLMs, even those enhanced with multimodal capabilities, offering a more reliable tool for marketers.
Therefore, we propose a system, called SOMONITOR, that combines predictive analytics and LLMs in a unified online marketing assistant. Importantly, marketing strategies encompass more than just data analysis; they involve understanding and implementing effective communication styles. It is essential to utilize these communication styles not merely as classification labels but as dynamic personas. This approach enables the identification of different communication schemas tailored to diverse user personas. These schemas, and the user personas themselves, can be analyzed in existing marketing campaigns via a clustering procedure (Section 4). The next step involves aligning these personas with appropriate communication styles to craft compelling narratives. These narratives, or stories, are essential content that marketers can directly employ in their campaigns. However, to construct these stories one has to use high-quality data, including marketing communications and advertisements that have previously proven to be successful in building rapport with the corresponding user personas.
Thus, the core issue circles back to the need for an advanced predictive model for content scoring, presented in Section 5. This model can be used to filter and identify “good” advertisements that can then serve to enhance the efficiency of LLMs in generating impactful marketing content. Our research addresses this need by introducing a predictive model that refines data input for LLMs, ensuring the generation of content that resonates well with target audiences and supports strategic marketing initiatives; in Section 6, we show how this model fits together with an LLM in a unified workflow that can generate complete user stories. The overall pipeline of SOMONITOR is illustrated in Fig. 1.
The paper is organized as follows: Section 2 surveys related work, in Section 3 we discuss LLM-based extraction of content pillars, Section 4 introduces persona and theme identification, Section 5 discusses content scoring, Section 6 introduces the story generation process, and Section 7 concludes the paper.
2. Related work
The all-encompassing success of LLMs has made them into a tool of choice for many applications (Makridakis et al., 2023; Minaee et al., 2024). Modern LLMs are able to perform a wide variety of tasks, from low-resource machine translation (Enis and Hopkins, 2024) to writing source code for full-scale applications (et al., 2021, 2023) and solving problems that require both mathematics and engineering design (Kevian et al., 2024). Importantly, there has been a growing body of research devoted to overcoming the quadratic complexity bottleneck inherent in the original Transformer architecture (Vaswani et al., 2017). Approaches include sparse attention mechanisms (Beltagy et al., 2020; Child et al., 2019), low-rank decomposition of attention weight matrices (Wang et al., 2020; Choromanski et al., 2022), and various modifications that combine self-attention with recurrent architectures which have inherently linear complexity (Ma et al., 2023; Orvieto et al., 2023). As a result, latest LLMs can extend their context windows (i.e., input sizes) up to millions of tokens (Team, 2024; Anthropic, 2024), which makes it possible to use LLMs directly for mid- and large-scale analysis of text datasets.
Recently, large language models have also obtained multimodal capabilities. They have been successfully used for video understanding (Tang et al., 2024), and the recently released GPT-4o, where ’o’ stands for “Omni”, brings seamless multimodality of image, audio, and video analysis to state-of-the-art LLMs (OpenAI, 2024). Moreover, LLMs can be combined with diffusion-based generative models (Ho et al., 2020; Song et al., 2021; Rombach et al., 2022) to achieve state-of-the-art image and video generation capabilities (Ho et al., 2022; Hong et al., 2022; Liu et al., 2024).
Being able to interpret data through human-perceivable dynamic categories, such as topics or audiences, represents a significant step towards explainable AI. Nonetheless, the sheer volume of content and its prioritization continue to pose challenges, with limitations inherent in LLMs, in particular related to their hallucination phenomena (Zhang et al., 2023). For instance, if a marketer understands that the primary audience for 1000 pieces of marketing content is “families”, but lacks data on the content’s popularity across this audience, such insights are not actionable for making informed decisions. At the same time, given a ranking task, LLMs always come up with different responses in consequent runs (Chen et al., 2023). At the same time, platforms dedicated to online content dissemination, such as social networking and search engine platforms, namely Facebook (Meta(He et al., 2014), Google (McMahan et al., 2013), Alibaba (Ouyang et al., 2019), and Taobao (Ge et al., 2018), have invested considerable efforts towards content prioritization and scoring. Unfortunately, these advancements are based on private datasets and are typically inaccessible to external content creators, including digital marketers. Proprietary engagement data, crucial for refining content scoring models, remains exclusive to these platforms, thereby hindering independent creators from accurately predicting or enhancing the efficacy of their content.
From all the above, it is evident that the comprehensive capabilities of LLMs position them as high-level assistants across a myriad of tasks within the realm of online marketing. However, there exist relatively few research papers devoted to these potential applications (Feizi et al., 2024). The marketing industry is currently only starting to incorporate LLMs into the workflows, with many recent whitepapers and business surveys of possible applications, problems, and potential breadth of adoption of LLMs (Thackray and Northin, 2023; Keni and ElBermawy, 2024). Our recent work (Yang et al., 2023) presented a comprehensive online marketing analysis system based on large language models; in this paper, we build upon this work and combine and further improve LLM-based analysis with predictive analytics. A key advantage noted in (Yang et al., 2023) was the fact that LLM answers are understandable for humans and actionable in terms of business results: LLMs such as GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023) were already able to explain the “reasoning” behind recommender models and provide aggregate insights about advertising campaigns. Before LLMs, text corpora in the context of recommender systems had been aggregated via topic modeling (Koltcov et al., 2014; Nikolenko, 2015), sometimes also with deep learning (Tutubalina and Nikolenko, 2018) and user profiling (Buraya et al., 2018; Tutubalina and Nikolenko, 2017), but topic modeling is based on the bag-of-words assumption and cannot summarize text as an LLM does. Visual understanding of ads has also been explored with convolutional networks (Savchenko et al., 2020).
However, a robust technology that enables effective differentiation of content sources for insight extraction, as well as a cohesive framework for data exploration that integrates both content differentiation and LLM-based explanations, is still sorely needed. Our work aims to address these gaps by developing methods that enhance the functionality of LLMs in online marketing applications, facilitating more efficient utilization of modern AI in real-world scenarios.
3. Content pillar identification
As noted above, data handlers in performance marketing frequently encounter large volumes of creative assets to analyze, along with numerous ongoing promotional activities from competitors in an “always-on” environment. Given this context, conducting detailed analyses for each advertisement or even for every specific competitor campaign becomes impractical, highlighting the need for automated analytical tools. These tools are crucial for assisting marketers in quickly assimilating, synthesizing, and driving inferences from extensive data and information available to them.
To address this challenge, we propose to integrate a novel method into our framework that leverages large language models (LLMs) to provide semantic insights into content, referred to as “content pillars” below. We outline an analytical pipeline that incorporates explanations and generations based on LLMs, demonstrating its practical utility through an empirical scenario involving four Singapore-based telecommunication companies. This component of our framework aims to enhance the interpretability and understanding of data, thereby supporting more informed decision-making in these dynamic and competitive sectors.
The architecture of our analytic process is illustrated in Fig. 2. Initially, we employ an LLM to extract specific, well-defined insights from the input content, such as the customer needs addressed by the advertisement and the products being advertised. These insights are then archived as content pillars in a tabular format. Subsequently, these pillars are utilized alongside custom-crafted prompts to perform a generalized analysis of a brand’s target demographics, personas, needs, and insights conveyed by the advertisements, as well as the tone and topical categories of both current and previous campaigns. This comprehensive analysis mirrors the campaign assessments typically conducted by marketing professionals and can be further used to refine a brand’s messaging, tone, and target demographics. Below we explore these elements in detail.
Figure 3 presents sample qualitative results of our experiments on the analysis of content pieces. We curated batches of advertisements from the Facebook Ad Library related to the same brand and processed them through an LLM. The processing was uniquely tailored through natural language prompt engineering. As a result, the LLM successfully identified key attributes of each advertisement, providing adept responses to ostensibly “human” inquiries such as pinpointing the human need, insight, and predominant archetypes in a specific ad. Notably, responses to most queries are standardized (as directed by the prompt) and amenable to automated processing. This type of analysis has always been pivotal in online marketing, and to our knowledge, it has not been effectively automated and scaled until now, traditionally requiring human intervention and thus being limited to a select number of advertisements rather than an extensive dataset.
Figure 2. General pipeline of our LLM-based analysis
Figure 3. Sample ad analysis
4. Communication Persona and Theme Identification
4.1. Problem setting
In the field of marketing, utilizing explainable data descriptors such as the content pillars is essential. However, to significantly enhance a brand’s marketing and communication strategies it is necessary to go beyond mere description and engage in inferential data analysis. Unfortunately, marketers often lack the necessary skills in data operations, hindering their ability to derive meaningful conclusions from the data, thereby requiring additional support. To bridge this gap, we propose a unified framework SOMONITOR that unites LLM-based analysis with content scoring.
SOMONITOR is designed to provide marketers with higher-level content observations.
In marketing, the narrative constructed by any content emerges from an underlying story, which is a synthesis of a communication theme and a specific customer persona that the campaign aims to engage. We will come back to the stories in Section 6, illustrating this concept through a representative example, highlighting the integral relationship between the storytelling elements and the targeted persona. This approach empowers marketers to craft more resonant and effective communication strategies by aligning their messaging with the expectations and needs of their intended audiences. But to get to a good story, we need to extract two essential storytelling components: personas and challenges.
4.2. Persona Identification
In marketing, the customer persona is best developed based on data about the audience of a brand, supported by other content pillars such as customer needs, interests, and aspirations (Malik, 2019). We have employed a commonly-used clustering algorithm X-Means (Pelleg et al., 2000) to cluster the content pieces into distinct customer personas based on the content from the brands, identifying the number of clusters automatically via Bayesian Information Criterion values (Pelleg et al., 2000). Clustering has been performed over the ADA embeddings, a state of the art representation learning model for text (Neelakantan et al., 2022), of the “Audience” content pillar as the main semantic component differentiating the data. X-Means was run with $3$ starting centroids and $k_{\max}=50$.
For this study, we analyzed data from the Gojek car-hailing brand and its main competitor in Singapore, Grab, focusing exclusively on the B2B business audience to limit the number of personas. The data for this experiment, collected over 2023 and 2024, comprised $5{,}967$ content pieces, with $3{,}703$ published as ads and $2{,}264$ as organic social media posts. Since most of the content related to B2B product offerings for both brands was presented in ads, we focused solely on the advertising content, totaling $1{,}120$ ads related to business products. Brand-wise, Gojek ran the majority of the ads ($76.8$%, or $849$ ads), while Grab accounted for the remaining $23.2$% ($271$ ads).
Figure 4. Sample persona analysis results
After detecting each cluster, we annotated them and generated parameters such as cluster name and cluster description by prompting an LLM to do so based on the predominant parameters and content of the posts associated with each cluster. Finally, an image for each persona was generated by the LLM in accordance with the persona’s description. As shown in Fig. 4, we identified three distinct clusters, each representing a unique customer persona along with their associated characteristics: “Efficiency Enthusiasts” (Persona 1), “Financial Empowerment Champions” (Persona 2), and “Efficiency Innovators” (Persona 3), with content counts of $206$, $144$, and $707$ respectively. Note that some content items did not correspond to any of the three identified personas (i.e., represented outliers) and were thus excluded from the final tally, explaining the disparity with the initial $1{,}120$ ads.
As a result, notable distinctions between the personas have emerged. Both Persona 1 and Persona 3 focus on organizational efficiency, albeit through different approaches: Persona 1 advocates for efficiency through traditional methodologies, while Persona 3 champions it through innovative strategies. In contrast, Persona 2 stands apart by primarily emphasizing the need for financial empowerment of employees.
4.3. Communication Theme: Challenges
A persona is an essential but not sufficient component needed to come up with a marketing story and its corresponding content briefs. In particular, it is important to identify a communication scheme to be combined with the persona’s key attributes such as background or character, so that the story would get proper context and would grab attention (Malik, 2019). To explore this, we have made a further step towards clustering our data, but this time we have performed X-Means clustering based on the “Insights” content pillar, again using ADA embeddings for text (Neelakantan et al., 2022) and running X-Means with $3$ starting centroids and $k_{\max}=50$.
Figure 5. Sample challenges mined by our system
Similarly to persona clustering, we annotated the resulting challenge clusters and generated parameters such as cluster name and cluster description by prompting an LLM to do so based on predominant parameters and content of the posts associated with each cluster. Figure 5 shows sample results: in this case, we have identified three distinct clusters, each representing a unique customer challenge along with their associated characteristics: “Inefficient Corporate Expense and Transportation Management” (Challenge 1), “Improving Business Operations and Employee Satisfaction” (Challenge 2), and “Streamlining Work Transport Processes” (Challenge 3), with content counts of $672$, $94$, and $336$ respectively.
5. Content scoring
5.1. Problem setting and dataset
To render the recommendations generated by the system actionable, one has to enable marketers to evaluate data and select which content to incorporate into analytics, ideally based on anticipated performance. This leads to the need for content scoring, so in this section we add to SOMONITOR a system for the scoring of advertising content, developed in our previous work (Yang et al., 2023), and evaluate it against state of the art LLMs to prove its continued relevance.
Brand nDCG@5 nDCG@10 Recall@3 Recall@5
A 0.498 0.796 0.333 0.4
B 0.541 0.821 0.333 0.4
C 0.588 0.829 0.6 0.6
Table 1. Experimental results for SOMONITOR.
nDCG@5 nDCG@10 Recall@3 Recall@5
GPT4-gd 0.591 0.818 0.333 0.6
GPT4 0.465 0.755 0 0.4
GPT4o-gd 0.599 0.827 0.333 0.6
GPT4o 0.511 0.750 0.333 0.4
Gemini-gd 0.578 0.817 0.333 0.6
Gemini 0.564 0.769 0 0.6
SOMONITOR 0.588 0.829 0.6 0.6
Table 2. Performance evaluation against LLM baselines; “gd” represents a grounded model.
To assess the effectiveness of our recommender system, we conducted a validation study using real ad creatives and their performance data. We collected real-world ad creatives and their performance data from three brands on Facebook, each with distinct marketing objectives: sales, conversion, and traffic, spanning from October 1, 2023 to December 31, 2023. The dataset comprised text and images of their ads alongside their performance measured as CTR values. We chose three (anonymized) brands: Brand A (16 ads in the fitness industry), Brand B (29 ads in the education industry), and Brand C (16 ads in the automobile industry).
The ads were assessed not just according to the brand affiliation, but also according to the Ad objective selected on the Meta Ads platform, which is a crucial component of evaluation in our scenario due to the fundamental differences in click-through rates across various industries and marketing goals that Ads achieve with different Ad objective settings (Yang et al., 2023). For instance, advertisements optimized for clicks may result in lower costs per click and higher click-through rates, yet they may not effectively convert customers. Conversely, ads aimed at achieving better conversion rates might not exhibit as high click-through rates as those optimized merely for clicks.
5.2. Content scoring models
For content scoring, we use our previously developed approach (Yang et al., 2023). Although this model was originally designed for predicting CTR by classifying advertisements into CTR performance categories, in this work we operate in a recommendation problem setting, i.e., pose content scoring as a ranking problem. This is because, first, marketers do not necessarily need precise CTR values to choose the best content but rather a comparative assessment, second, CTR values undergo significant distribution shift across platforms, time, and other external factors, but the ranking of ads should be more stable, and third, the LLM performance in qualitative ranking should be significantly better than in numerical predictions, ensuring a more fair comparison.
Thus, we take the SODA model from (Yang et al., 2023) that predicts a distribution over “High CTR”, “Average CTR”, and “Low CTR” labels and add a linear layer that translates the probabilities into a single score via a trainable linear combination:
$\mathrm{Score}=\alpha\cdot p(\mathrm{HighCTR})+\beta\cdot(1-p(\mathrm{HighCTR})).$
As a result, we convert a pretrained classification model into a ranking model.
On the side of LLMs, we selected state of the art models GPT-4, GPT-4o, and Gemini 1.5 Pro for comparison with SOMONITOR. GPT-4 and its optimized version, GPT-4o, represent the latest advancements in the GPT series, known for their robustness and versatility in handling diverse text-based and multimodal tasks (OpenAI, 2023), while Gemini 1.5 Pro is a powerful model developed by Google and competing with OpenAI’s models successfully in many benchmarks (Team, 2024).
However, a significant challenge in utilizing LLMs for this task is their tendency to hallucinate, which can result in generating inaccurate or nonsensical information that compromises the reliability of their output. To mitigate this issue and ground the LLM-based approach to ensure the predicted data was trustworthy, we implemented several measures to reduce randomness:
  • temperature adjustment: we set the temperature parameter of the LLMs to 0.1, which is known to improve stability by reducing the variability in the model’s outputs;
  • grounding: to further mitigate randomness, we included another three example advertisements from the brand within the prompt, each with a specified CTR rank: one with the best performance, one with average performance, and one with the worst performance;
  • LLM Ensemble: inspired by ensemble methods, we conducted the ranking process with LLMs five times using the same settings, and the final rankings were determined by summing the rank positions of each advertisement from these five runs and then ordering them from the lowest to highest total rank.
5.3. Results
For evaluation, we use standard ranking-based metrics. Specifically, we measured Recall@3 and Recall@5 to gauge the model’s ability to retrieve relevant items within the top 3 and top 5 recommendations; the recall metric is very important in practice, especially in resource-constrained domains such as online advertising campaigns. High Recall@$k$ ensures that top $k$ recommendations encompass a substantial portion of the truly relevant items, which enables marketers operating under tight budgets to focus their efforts on the most promising opportunities while minimizing the risk of overlooking potentially valuable prospects. We have also computed normalized discounted cumulative gain (nDCG) at cutoffs 5 and 10, another commonly used way to assess ranking quality.
First, we present a case study in Table 1 that showcases the framework’s performance across three major industries. In the fitness sector, SOMONITOR achieved moderate effectiveness with nDCG@5 of 0.498 and nDCG@10 of 0.796, alongside Recall@3 and Recall@5 scores of 0.333 and 0.4, respectively. Brand B, representing education, showed improved performance, with nDCG@5 of 0.541 and nDCG@10 of 0.821, but recall stayed the same. In the automobile industry results improved, achieving nDCG@5 of 0.588 and nDCG@10 of 0.829, with recall@3 and recall@5 both scoring 0.6. These findings show that content scoring quality varies across industries and highlight the importance of industry-specific optimization to enhance recommendation accuracy.
Then we undertook a comparative evaluation of our SOMONITOR framework against LLM baselines. In campaigns focused on driving traffic, the primary goal is to maximize click-through rates (CTR), which serve as a measure of user engagement and interaction with the advertisement, which is particularly pertinent given the distinctive challenges posed by traffic-driven campaigns as opposed to conversion-oriented strategies. Thus, we evaluate the performance with the data from Brand C, a traffic-oriented campaign. In Table 2, we present the evaluation result from a case study focused on Brand C, comparing the performance of SOMONITOR against modern large language models.
A key observation from Table 2 is the significant performance boost achieved by incorporating knowledge grounding across various models and metrics. For instance, the nDCG@5 score for GPT-4 improved from 0.465 without grounding to 0.591 with grounding, while its Recall@5 increased from 0.4 to 0.6. This trend is consistent across other LLM baselines, such as GPT-4o and Gemini1.5 Pro, underscoring the critical role of grounding in enhancing the relevance and quality of recommendations. Ungrounded LLMs exhibit markedly lower performance, particularly in the Recall@3 metric: both Gemini and GPT-4 failed to place any high-performance items within their top $3$ results at all.
As a result, the prediction model outperforms LLMs in recommending relevant items, achieving the highest nDCG@10 of 0.829 and recall@3 of 0.6 among the compared models. This superior performance demonstrates SOMONITOR’s ability to effectively balance ranking accuracy and item recall, delivering highly relevant and stable (especially compared to individual LLM results) recommendations to users. Notably, these outcomes are achieved without any fine-tuning (grounding) using data from the actual industry-specific distributions: we used a pretrained model. This offers significant potential for further improvement of results should such fine-tuning be implemented, a direction we intend to explore in our future work. Improvements in LLMs’ Recall@k due to grounding also hint at the critical role of domain knowledge and context in enhancing the relevance and quality of recommendations. Ungrounded LLMs, relying solely on their pre-trained language models, struggle to capture the nuances and specifics of the recommendation task, resulting in sub-optimal item retrieval. By incorporating domain-specific knowledge and grounding the models in the relevant context, the recommendations become more tailored and relevant.
We note that our proposed SOMONITOR framework outperforms the baselines in both Recall@3 and Recall@5, achieving a Recall@3 score of 0.6 which is higher than any LLM’s result. This supports our conclusion that the online marketing industry still cannot fully rely on LLMs for content scoring and explains our decision to include a CTR prediction and ranking model into the mostly LLM-based SOMONITOR framework.
6. Data-Driven Story Generation
Figure 6. User story generation: pipeline and an example.
In order to establish a data-driven narrative that would be useful, easy to understand, and attractive to content marketing professionals, it is essential to delineate the content and marketing strategies a brand may deploy, guided by the analysis of the best-performing content from competitive brands. At this point, we have already discussed how the SOMONITOR framework extracts personas and challenges based on advertising content prefiltered by a CTR prediction and ranking model. Now it is time to combine them into a unified marketing story.
Consider the scenario where we are part of a marketing team at Gojek, tasked with identifying a suitable communication theme and a target persona for a campaign aimed at market expansion through new audience engagement. It is desirable to opt for a persona and theme that are presently underexploited, thereby expanding our audience reach. As illustrated in Fig. 4, the analysis shows that Persona 1, labeled “Efficiency Enthusiasts”, is more heavily leveraged by Grab than by Gojek, indicating a potential niche opportunity for Gojek. At the same time, as shown in Fig. 5, the communication theme “Challenge 3: Streamlining Work Transport Processes” is more frequently addressed by our competitor, Grab, but is underrepresented in Gojek’s content. This observation advocates for adopting this theme for persona-centric narrative generation.
Consequently, adhering to the framework outlined in Figure 6, we prompt an LLM to amalgamate the characteristics of the selected persona and the chosen communication theme to devise a narrative for a fictitious character, “Samuel Tan”, a business owner in Singapore. The pipeline for this generation is shown in Figure 6: character description is also generated by a custom-prompted LLM, and then the character and theme are combined to create a compelling story. The story is abbreviated in Fig. 6, in reality it explains over several paragraphs how Samuel prioritizes business efficiency but often neglects employee satisfaction, leading to increased turnover rates. Recognizing this issue, Samuel opts for Gojek not merely as an efficiency solution but as a means to reengage with employees, enhance their efficiency, and improve overall job satisfaction; the story ends with a conclusion (emphasized in Fig. 6) that shows an important insight that Samuel has obtained by engaging with Gojek. This narrative could serve as a content brief for a creative agency or be utilized as a pitch for a novel marketing campaign, previously unexplored by the marketing team in such depth and with such rapid production capability. In our professional opinion, content briefs produced by SOMONITOR are on par with similar human-written narratives and can be used just as effectively.
7. Conclusion
In this work, we have introduced an innovative explainable AI framework SOMONITOR that aims to combine human intuition with the efficiency of modern AI to significantly enhance the capabilities of marketers across all stages of the marketing funnel, from strategic planning to content creation and campaign execution. The challenges traditionally faced by marketers, especially the need to manually sift through immense quantities of transient online content, are addressed by SOMONITOR’s ability to automate and refine the process of competitor analysis, content research, and strategic branding. SOMONITOR ranks advertising content by its potential to achieve higher click-through rates via a specially trained model and then employs LLMs to distill core content pillars such as target audiences, customer needs, and product features from successful competitor marketing endeavors. These pillars then lead to communication themes and targeted customer personas, providing a structured basis for content and campaign planning. By integrating these insights with data from a brand’s own marketing campaigns, SOMONITOR not only constructs compelling narratives for new customer personas but also generates detailed content briefs in the form of customer persona stories, empowering marketing teams to produce content and execute campaigns. We believe that SOMONITOR can transform digital marketing operations, enabling marketers to quickly navigate through large datasets and derive actionable insights that boost campaign effectiveness and enhance job satisfaction, redefining traditional marketing approaches (Farseev, 2024).